# Large Language Model-Driven Closed-Loop UAV Operation with Semantic Observations

### v1

@@ Abstract
Large Language Models (LLMs) have revolutionized robotic autonomy, including Unmanned Aerial Vehicles (UAVs). Recent studies have demonstrated the potential of LLMs for translating human instructions into executable control code for UAV operations. However, LLMs still face challenges from logical reasoning and complex decision-making, leading to concerns about the reliability of LLM-driven UAV operations. In this paper, we propose a LLM-driven closed-loop control framework that enables reliable UAV operations powered by effective feedback and refinement using two LLM modules, i.e., a Code Generator and an Evaluator. Our framework transforms numerical state observations from UAV operations into natural language trajectory descriptions to enhance the evaluator LLM’s understanding of UAV dynamics for precise feedback generation. Our framework also enables a simulation-based refinement process, and hence eliminates the risks to physical UAVs caused by incorrect code execution during the refinement. Extensive experiments on UAV control tasks with different complexities are conducted. The experimental results show that our framework can achieve reliable UAV operations using LLMs, which significantly outperforms baseline approaches in terms of success rate and completeness with the increase of task complexity.
@@
UAV, Large Language Model, Robotics.
@@ I Introduction
Uavs, or drones, are increasingly being integrated into the Internet of Things (IoTs) ecosystems to support various consumer applications, such as precision agriculture, infrastructure inspection, and surveillance [1, 2, 3, 4]. Their high mobility and rich sensing capabilities make them an ideal candidate for data collection, communication, and service delivery tasks in these applications. Designing UAVs that have intelligence to perform complex tasks in these applications correctly remains challenging, as it demands both robust instruction interpretation and executing tasks with robust reasoning. Recent advances in LLMs [5, 6, 7, 8] have demonstrated remarkable proficiency in robotic areas such as planning [9], navigation [10, 11], communication [12], and automation [13]. With LLMs’ exceptional natural language (NL) understanding and context generation capabilities, LLMs offer a promising avenue for robots with greater intelligence. Motivated by these developments, research efforts have been made towards the reliable integration of LLMs and UAVs [14, 15, 16, 17, 18], to enable LLM-driven UAV operations.
@@ II-B Prompt Engineering
Recent studies have increasingly employed prompt engineering [20] to improve the reliability of LLM-driven robotic systems. For instance, GSCE proposes a prompt framework as a system prompt [33] to improve the reliability of LLM-driven UAV control code generation [18]. Prompt engineering typically leverages two core techniques: in-context learning [34] and Chain-of-Thought (CoT) reasoning [35]. In-context learning enables the LLM to identify patterns from a limited set of examples and apply them to new situations. This facilitates the generation of control code that adheres to robot policy and how to ground task descriptions from a few examples [21, 36]. On the other hand, CoT prompts the LLM to articulate intermediate inference steps, which is valuable for robotic tasks requiring sequential, stepwise decision-making. By decomposing a task into steps, CoT encourages the production of code that aligns with each stage of the intended action plan. Prior studies have embedded CoT within the examples to guide LLMs through reasoning step-by-step via in-context learning [24, 36]. In this study, we integrate both CoT and in-context learning strategies to enhance the quality of code generation and evaluation in LLM-driven UAV control.

### v2

@@ Abstract
Recent advances in large Language Models (LLMs) have revolutionized mobile robots, including unmanned aerial vehicles (UAVs), enabling their intelligent operation within Internet of Things (IoT) ecosystems. However, LLMs still face challenges from logical reasoning and complex decision-making, leading to concerns about the reliability of LLM-driven UAV operations in IoT applications. In this paper, we propose a LLM-driven closed-loop control framework that enables reliable UAV operations powered by effective feedback and refinement using two LLM modules, i.e., a Code Generator and an Evaluator. Our framework transforms numerical state observations from UAV operations into natural language trajectory descriptions to enhance the evaluator LLM’s understanding of UAV dynamics for precise feedback generation. Our framework also enables a simulation-based refinement process, and hence eliminates the risks to physical UAVs caused by incorrect code execution during the refinement. Extensive experiments on UAV control tasks with different complexities are conducted. The experimental results show that our framework can achieve reliable UAV operations using LLMs, which significantly outperforms baseline approaches in terms of success rate and completeness with the increase of task complexity.
@@
UAV, Large Language Model, Robotics, IoT.
@@ I Introduction
Uavs, or drones, are increasingly being integrated into IoT ecosystems to support various applications, such as surveillance and monitoring, smart cities, and infrastructure inspection [1]. Their high mobility and rich sensing capabilities make them an ideal candidate for data collection, communication, and service delivery tasks in these applications. Designing UAVs that have intelligence to perform complex tasks in these applications correctly remains challenging, as it demands both robust instruction interpretation and executing tasks with robust reasoning. Recent advances in LLMs [2, 3, 4, 5] have demonstrated remarkable proficiency in robotic areas such as planning [6], navigation [7, 8], communication [9], and automation [10]. With LLMs’ exceptional natural language (NL) understanding and context generation capabilities, LLMs offer a promising avenue for robots with greater intelligence. Motivated by these developments, research efforts have been made towards the reliable integration of LLMs and UAVs [11, 12, 13, 14, 15], to enable LLM-driven UAV operations.
@@ II-B Prompt Engineering
Recent studies have increasingly employed prompt engineering [17] to improve the reliability of LLM-driven robotic systems. For instance, GSCE proposes a prompt framework as a system prompt to improve the reliability of LLM-driven UAV control code generation [15]. Prompt engineering typically leverages two core techniques: in-context learning [30] and Chain-of-Thought (CoT) reasoning [31]. In-context learning enables the LLM to identify patterns from a limited set of examples and apply them to new situations. This facilitates the generation of control code that adheres to robot policy and how to ground task descriptions from a few examples [18, 32]. On the other hand, CoT prompts the LLM to articulate intermediate inference steps, which is valuable for robotic tasks requiring sequential, stepwise decision-making. By decomposing a task into steps, CoT encourages the production of code that aligns with each stage of the intended action plan. Prior studies have embedded CoT within the examples to guide LLMs through reasoning step-by-step via in-context learning [21, 32]. In this study, we integrate both CoT and in-context learning strategies to enhance the quality of code generation and evaluation in LLM-driven UAV control.